### v1

# Quantum speed limit in quantum sensing

Abstract
Quantum sensors capitalize on advanced control sequences for maximizing sensitivity and precision. However, protocols are not usually optimized for temporal resolution. Here, we establish the limits for time-resolved sensing of dynamical signals using qubit probes. We show that the best possible time resolution is closely related to the quantum speed limit (QSL), which describes the minimum time needed to transform between basis states. We further show that a bipartite control sequence consisting of two phase-shifted pulses reaches the QSL. Practical implementation is discussed based on the example of the spin-1 qutrit of a nitrogen-vacancy (NV) center in diamond.
The energy-time uncertainty principle introduced by Heisenberg is a fundamental concept of quantum mechanics. While formulated loosely in Heisenberg’s initial work, Robertson [1] and Bohr [2] put it on firm ground by formalizing the relationship between uncertainty and non-commutativity of observables. Twenty years later, Mandelstam and Tamm [3] showed that the energy-time uncertainty is not an uncertainty relation due to non-commutativity, but rather a statement about intrinsic time scales of quantum systems [4]. This insight led to the derivation of a quantum speed limit (QSL). To do so, Mandelstam and Tamm used the von Neumann equation with the projection operator to develop an expression for the overlap between time-evolved states,
$\braket{\psi(0)}{\psi(t)}\geq\cos\left(\frac{\langle\Delta H\rangle t}{\hbar}\right)\ ,$ (1)
where $\langle\Delta H\rangle^{2}=\langle H^{2}\rangle-\langle H\rangle^{2}$ is the variance of the Hamiltonian. The minimum time needed to obtain a fully orthogonal state, i.e., $\braket{\psi(0)}{\psi(t)}=0$, is then given by
$t\geq t_{\mathrm{QSL}}^{(\mathrm{MT})}=\frac{\pi}{2}\frac{\hbar}{\langle\Delta H\rangle}\ .$ (2)
Later, Margolus and Levitin [5] proposed an alternative route to deriving a QSL based on the integrability of the Schrödinger equation to obtain a maximum evolution speed, resulting in
$t\geq t_{\mathrm{QSL}}^{(\mathrm{ML})}=\frac{\pi}{2}\frac{\hbar}{\langle H\rangle}\ .$ (3)
In contrast to the Mandelstam-Tamm definition, Eq. (3) bounds the time on the mean energy $\langle H\rangle$ defined relative to the energy of the ground state. For a two-level system, the two QSLs coincide while for systems open to a continuum of states or for systems consisting of more than two states, such as qutrits [6], the unified bound is tight [7, 8].
Since the initial formulations of the QSL, theoretical work has focused on more complex situations such as open quantum systems [9] and systems with quantum entanglement [10], as well as applications in quantum information processing [11]. Here, the QSL sets the maximum speed at which computations can be performed and therefore permits deriving an upper bound of the computational limits of the universe [12, 13]. Further, the QSL has been studied in the field of quantum optimal control [14], quantum thermodynamics [15], and metrology with respect to quantum clocks [16]. Obviously, the finite time of qubit operations will also limit the time resolution achievable in quantum sensing tasks [17]. The time limitation is of practical importance because of the many existing and envisioned applications of quantum sensors.
In this work, we investigate the relation between the QSL and the time resolution achievable in quantum sensing experiments. We show that a bipartite sensing sequence consisting of two phase-shifted control rotations, equivalent to a Ramsey sequence with zero time delay, reaches the QSL. Opposite to a Ramsey measurement, however, phase accumulation occurs during control rotations rather than a free evolution interval. We derive quantitative expressions for the quantum phase pick-up as a function of control rotation angle and velocity. We use these expressions to define the time resolution and bandwidth of the sensing sequence. We also show that time resolution can be extended beyond the QSL by trading for a reduced signal-to-noise ratio. As an example, we simulate the expected response for a nitrogen-vacancy (NV) center in diamond that is exposed to rapidly varying magnetic signal.
Figure 1: Concept of time-resolved sensing at the QSL. (a) Canonical control sequence for signal estimation via quantum phase accumulation. $\hat{R}$ and $\hat{R}^{\prime}$ are control rotations, and $\hat{R}^{\phi}$ is the qubit rotation due to interaction with the signal. (b) Stopwatch analogy of (a). (c) Pulse-timing diagram of the control sequence for infinitely fast control rotations. Phase accumulation occurs entirely during the free evolution time $\tau$. $\pi/2$ are rotation angles and $x$, $-y$ are axes of rotation. (d) Pulse-timing diagram for finite duration $t_{\mathrm{R}}>0$ of control rotations. (e) Pulse-timing diagram at the QSL, when the interpulse delay becomes zero and $\tau=2t_{\mathrm{R}}$. Phase accumulation now occurs entirely during control rotations. $\alpha$ is the rotation angle. (f) Bloch sphere trajectories for sequence (e) with $\alpha=90^{\circ}$. Dots are the projection on the $Z$ axis. Trajectories are shown for $\phi=0^{\circ}$ (light blue), $10^{\circ}$, $20^{\circ}$ and $30^{\circ}$ (dark blue). (g), Bloch sphere trajectories for $\alpha=63^{\circ}$. Further trajectories are shown in [18].
The canonical quantum sensing scheme of Ramsey interferometry is shown in Fig. 1(a) [19, 17]. In its most basic form, the scheme uses two state transformations to initiate and halt the coherent evolution of a probe qubit subject to an external signal. The first transformation rotates the qubit from a known initial basis state |$\ket{\psi(0)}=\ket{0}$ into a coherent superposition state $\frac{1}{\sqrt{2}}(\ket{0}+\ket{1})$, which then evolves for a given time $\tau$ into state $\frac{1}{\sqrt{2}}(\ket{0}+e^{-\mathrm{i}\phi(\tau)}\ket{1})$, thereby acquiring a quantum phase
$\phi=\int_{0}^{\tau}[\omega_{0}+\delta\omega(t^{\prime})]\,\mathrm{d}t^{\prime}\ .$ (4)
Here, $\hbar\omega_{0}$ is the static energy gap between the qubit’s energy levels, and $\hbar\delta\omega(t)$ accounts for a small, time-dependent modulation due to the presence of the signal. After coherent evolution, a second state transformation rotates the qubit back to the original basis, followed by projective state readout  [^1] In general, the initial and readout bases may be different, but for simplicity and without loss of generality, we assume them to be the same [17]. By this, the canonical quantum sensing experiment measures the expectation value of the projector onto the initial state |$\ket{0}$. The state transformations can be described by two control rotations, $\hat{R}$ and $\hat{R}^{\prime}$, respectively. These are analogous to the start and stop triggers in a classical test and measurement task (Fig. 1(b)). There, the time elapsed between start and stop events determines the time resolution.
Ideally, the control rotations $\hat{R}$ and $\hat{R}^{\prime}$ are infinitely fast [Fig. 1(c)]. However, owing to the QSL, $\hat{R}$ and $\hat{R}^{\prime}$ must have a finite duration $t_{\mathrm{R}}$ or, equivalently, a finite angular velocity $\Omega$. In experiments, the maximum $\Omega$ may be limited by many factors, including finite available driving power, competition between $\omega_{0}$ and $\Omega$, excitation of further lying energy levels, or a combination of those. The finite speed of control rotations fundamentally limits the time resolution of the sensing sequence.
For finite $t_{\mathrm{R}}$, qubit evolution under the signal Hamiltonian already occurs during the control rotations. The qubit-acquired phase then becomes the sum of the phase pick-up during the free evolution interval and the phase pick-up during control rotations [Fig. 1(d)]. In the most extreme case, the free evolution time is zero, and phase pick-up occurs entirely during control rotations [Fig. 1(e)]. In this situation, the time resolution of the sequence reaches that of the QSL.
We next derive quantitative relationships between the phase pick-up $\phi_{\mathrm{eff}}$ and the velocity $\Omega$ and angle $\alpha$ of control rotations. For our derivation, we focus on a bipartite control sequence consisting of two consecutive qubit rotations $\hat{R}=\hat{R}_{Y}^{\alpha}$ and $\hat{R}^{\prime}=\hat{R}_{\pm X}^{\alpha}$ of equal duration $t_{\mathrm{R}}$ [Fig. 1(e)]. Here, $X$ and $Y$ are orthogonal axes, as shown in Fig. 1(f). This sequence represents the basic Ramsey scheme with zero time delay. The sequence duration is $\tau=2t_{\mathrm{R}}=\alpha/\Omega$, where $\Omega$ is the maximum allowed rotation velocity. In the Supplemental Material [18], we show that an equal time-share ($t_{\mathrm{R}}=t_{\mathrm{R^{\prime}}}$) between the two parts of the bipartite pulse sequence yields optimum sensitivity (as defined by Eq. (8) below). This is also true when varying the angle between rotation axes.
Our sensor output is the overlap between the initial qubit state |$\ket{\psi(0)}=\ket{0}$ and the final qubit state |$\ket{\psi(\tau)}$, given by the transition probability [17, 10]
$p$ $=1-|\braket{0}{\psi(\tau)}|^{2}$ (5a)
$=1-|\bra{0}\hat{R}\hat{R}^{\prime}\ket{0}|^{2}$ (5b)
$=1-|\bra{0}\hat{R}_{\pm X}^{\alpha}\hat{R}_{Y}^{\alpha}\ket{0}|^{2}.$ (5c)
In the limit of strong control fields ($\Omega\gg\delta\omega(t)$), within linear response ($\phi\ll\pi/2$), and assuming that $\delta\omega(t)\approx\delta\omega$ is stationary during $\tau$, the transition probability of this sequence is [18]
$p$ $=\underbrace{\frac{1}{4}(1-\cos 2\alpha)}_{p_{0}}+\underbrace{\frac{1}{2}\frac{\sin\alpha(\cos\alpha-1)}{\alpha}\phi}_{\delta_{p}}\ .$ (6)
Here, $p_{0}$ is the bias point of the measurement, and $\delta p$ is the change in probability due to the presence of the signal $\delta\omega$[17]. Eq. (6) defines an effective phase
$\delta p=\frac{1}{2}\phi_{\mathrm{eff}}=\frac{1}{2}\frac{\sin\alpha(\cos\alpha-1)}{\alpha}\phi:=\frac{1}{2}\epsilon\phi$ (7)
that corresponds to the ideal Ramsey phase $\phi=\delta\omega\tau$ [Eq. (4)] reduced by the scaling factor $\epsilon<1$. Using Eq. (7), we can define the sensitivity $\eta$ of the sequence by
$\eta:=d[\delta p]/d[\delta\omega]\ ,$ (8)
taken in the limit $\delta\omega\rightarrow 0$.
For the canonical case of qubit rotations between orthogonal axes on the Bloch sphere ($\alpha=90^{\circ}$), the transition probability is
$p=\frac{1}{2}+\frac{\phi}{\pi}$ (9)
corresponding to $p_{0}=0.5$ and $\epsilon=2/\pi\approx 0.637$. Fig. 1(f,g) shows Bloch-sphere trajectories for $\alpha=90^{\circ}$ and $\alpha<90^{\circ}$, and Fig. 2 plots $\phi_{\mathrm{eff}}$ as a function of $\tau$.
Figure 2: Phase pick-up $\phi_{\mathrm{eff}}/\phi$ as a function of $\tau$ for control rotations with finite velocity. For durations shorter than the QSL ($\tau<2t_{\mathrm{R}}$, solid curve), $\phi_{\mathrm{eff}}$ is determined by the sequence of Fig. 1(e). For durations longer than the QSL ($\tau>2t_{\mathrm{R}}$, dashed curve) $\phi_{\mathrm{eff}}$ is determined by the sequence of Fig. 1(d). The upper bound ($\phi_{\mathrm{eff}}=\phi$, dotted line) is for the hypothetical case of infinitely fast control rotations, corresponding to the sequence in Fig. 1(c).
Figure 3: Qubit output and sensing kernel. (a) Timing diagram. To sample the signal transient $B(t)$, the control sequence (second line) is stepped along $t$ in increments of $\Delta t$. The output signal, given by state probability $p(t)$, is the convolution between $B(t)$ and the kernel $k(t)$ of the control sequence. Gray blocks represent initialization and readout of the qubit state, and white blocks represent the control rotations. (b) Kernel $k(t)$ for control rotation angles $\alpha=22.5^{\circ}$ (light blue), $45^{\circ}$, $67^{\circ}$ and $90^{\circ}$ (dark blue). Kernels are computed using a lab-frame simulation of the spin evolution [18]. The time resolution $t_{\mathrm{min}}$ is defined by the full width at half maximum. (c) Normalized Bode plots $K(\omega)$ for the kernels shown in (b). The bandwidth $\Omega_{\text{BW}}$ is defined by the first root of $K(\omega)$. (d) Sensitivity $\eta\propto K(\omega)$ as a function of signal frequency $\omega$ and pulse duration $\tau$ at fixed Rabi frequency $\Omega$. For dynamical signals $\omega>0$, an optimum pulse duration exists where sensitivity is maximized (white curve).
Having established the phase pick-up of the bipartite control sequence, we proceed to the problem of sampling a time-dependent signal. As a generic example, we consider the detection of a transient magnetic field $\vec{B}(t)$ using a spin-1/2 system as the qubit probe where $\delta\omega(t)=\gamma B(t)$. Here, $\gamma$ is the transduction factor (given by the gyromagnetic ratio), and $B(t)$ is the vector component of $\vec{B}(t)$ parallel to the quantization axis of the spin qubit. The influence of off-axis components of $\vec{B}(t)$ is discussed in [18].
To record a time transient, we sample $B(t)$ point-by-point by incrementing the delay time $t$ of the control sequence with respect to a common start trigger at $t=0$ [Fig. 3(a)]. For signals that vary only slowly with time, $B(t)$ is almost stationary during control pulses, and the sensor output $\delta p(t)=\epsilon\tau\gamma B(t)$ is directly proportional to the signal field. On the other hand, for more rapidly changing signals, the acquired quantum phase becomes a more complex function of signal and control fields. In a general case, we express the transition probability $\delta p$ by the convolution
$\delta p(t)=\int_{t^{\prime}=-\infty}^{\infty}k(t^{\prime}-t)\gamma B(t^{\prime})\text{d}t^{\prime}$ (10)
where $k(t)$ is the kernel of the control sequence and $t^{\prime}$ the time delay between signal and control fields [Fig. 3(a)]. Specifically, for the bipartite sequence of Fig. 1(e), the sensing kernel is
$k(t)=\begin{cases}\sin[\Omega(\tau/2-|t|)]&|t|<\tau/2\\ 0&|t|>\tau/2\end{cases}\ .$ (11)
In the frequency domain, the kernel is given by the transfer function $K(\omega)=|\mathrm{FT}[k(t)]|$,
$K(w)=\frac{\sqrt{\frac{2}{\pi}}\Omega\left|\cos\left(\Omega\tau/2\right)-\cos\left(\omega\tau/2\right)\right|}{\left|\Omega^{2}-\omega^{2}\right|}\ ,$ (12)
where FT is the Fourier transform.
Figs. 3(b,c) plots kernel profiles $k(t)$ and corresponding transfer functions $K(\omega)$ for several rotation angles $\alpha=\Omega\tau$ between $0^{\circ}$ and $90^{\circ}$. Clearly, a smaller $\alpha$ leads to narrower kernels and, thus, an improved time resolution. Defining the time resolution $t_{\mathrm{min}}$ by the full width at half maximum of $k(t)$, we find for the bipartite control pulse that
$t_{\mathrm{min}}:=\tau\left(1-\frac{\arcsin\frac{\sin\alpha}{2}}{\alpha}\right)$ (13)
(See [18] for other possible definitions of $t_{\mathrm{min}}$). Specifically, $t_{\mathrm{min}}=\frac{2}{3}\tau=\frac{2\pi}{3\Omega}$ for $\alpha=90^{\circ}$, and $t_{\mathrm{min}}\approx\frac{1}{2}\tau=\frac{\alpha}{2\Omega}$ in the limit $\alpha\rightarrow 0$. Further, we can define a frequency bandwidth by the first root of the transfer function $K(\omega)$, given by
$\Omega_{\text{BW}}:=\Omega\left(\frac{2\pi}{\alpha}-1\right),$ (14)
where $\Omega_{\text{BW}}=3\Omega$ for $\alpha=90^{\circ}$ (Fig. 3(c) and [18]).
For a given angular velocity $\Omega$, shorter rotation angles $\alpha<90^{\circ}$ therefore provide a means to further improve the time resolution. The improvement is approximately $t_{\mathrm{min}}\propto\alpha$, and correspondingly for the bandwidth, $\Omega_{\text{BW}}\propto\alpha^{-1}$. The improved time resolution, however, comes at the expense of a drastically lowered sensitivity, since $\eta\propto\alpha^{2}$ for small $\alpha$ [Eqs. (7,8)]. The sensitivity is further illustrated in Fig. 3(d), which plots $\eta$ as a function of $\tau$ for arbitrary signal frequencies $\omega$.
The apparent improvement of the time resolution beyond the QSL for short $\alpha$ [Eq. (13)] is in accordance with the interpretation of the uncertainty principle as defined by Eq. (1): the QSL reflects the minimum time required for transferring a quantum state to a fully orthogonal state. By reducing the orthogonality, thus not fully transferring the state, the time requirement shrinks therefore permitting a higher time resolution.
As a practical example, we consider the $S=1$ spin system of a single nitrogen-vacancy (NV) center in diamond [21]. The NV center is a prototypical qubit sensor with a growing range of applications in materials science, physics, chemistry, and biology [22, 23, 24, 25], including the study of dynamical excitations in these systems [26, 27, 28]. The NV center exhibits three spin energy levels $m_{S}=0$ and $m_{S}=\pm 1$ and two allowed spin transitions with frequencies $\omega$ and $\omega^{\prime}$, as depicted in Fig. 4(a). To form an effective two-level system, the $\omega$ (or $\omega^{\prime}$) transition can be isolated by applying an axial bias field $B_{0}$ along the NV symmetry axis [Fig. 4(b)]. For driving fields $\Omega$ smaller than the frequency difference $|\omega^{\prime}-\omega|$ between the $m_{S}=\pm 1$ levels [Fig. 4(c)], the excitation is selective and the other ($\omega^{\prime}$ or $\omega$) transition can be neglected.
Figure 4: Ultimate limits to time resolution for an NV probe qutrit. (a) NV spin energy levels and allowed transitions $\omega$ and $\omega^{\prime}$. (b) Scaling of transition frequencies $\omega$ and $\omega^{\prime}$ with axial bias field $B_{0}$. $D=2.87\,\mathrm{GHz}$ is the zero-field splitting parameter. (c) Isolation of transition frequencies $|\omega^{\prime}-\omega|$ as a function of $B_{0}$. The frequency difference saturates at $2D$. (d) Numerical simulation of the state probability $p$ as a function of driving frequency $\Omega$. $p_{0}$ is the bias point and $\delta p$ the probability change caused by the signal, see Eq. (6). At high $\Omega/2\pi>D$, the probability signal inverts and goes to zero for $\Omega\rightarrow\infty$ for preparation and measurement in the $m_{S}=0$ basis (blue), but not in the $m_{S}=\pm 1$ (green) basis. For the simulation, a fixed value of $\delta\phi=0.2$ was taken independent of $\Omega$.
Fig. 4(d) shows the expected output probability $p$ for an NV spin qubit as a function of $\Omega$. The output probability is calculated using a laboratory-frame simulation of the spin dynamics, taking the full $S=1$ nature of the NV center and the effect of counter-rotating terms in the excitation pulses into account [18]. Two simulations are presented where the spin is either initialized and read out in the $m_{S}=0$ state or in the $m_{S}=-1$ state, respectively.
First, we consider the $m_{S}=0$ case. For moderate driving fields $\Omega$ that are safely smaller than any of the transition frequencies ($\omega$, $\omega^{\prime}$, $|\omega^{\prime}-\omega|$), the output probability $p$ is nearly independent of $\Omega$ and follows the theory of Eq. (7). On the other hand, for larger $\Omega$, the output probability is modified due to strong driving effects – namely, Bloch-Siegert shifts and breakdown of the rotating-wave approximation [29, 30, 31] – and by spurious excitation of the $\omega^{\prime}$ transition. Both effects are not accounted for by the basic theory of Eq. (7). Further, because the $\omega$ and $\omega^{\prime}$ transitions lead to opposite signs in the phase $\phi_{\mathrm{eff}}$, the signal is completely canceled for very large $\Omega\gg|\omega^{\prime}-\omega|$ (blue trace in Fig. 4(d)).
By contrast, when preparing and reading out the $m_{S}=-1$ spin state (green trace in Fig. 4(d)), the cancellation remains incomplete at any bias field. This provides a remedy to the $m_{S}=0$ case. The incomplete cancellation is due to selection rules for qubit rotations: transitions are allowed between $m_{S}=0$ and $m_{S}=\pm 1$ but forbidden between $m_{S}=-1$ and $m_{S}=+1$ states. Preparation and readout of $m_{S}=\pm 1$ are realized experimentally by slow, selective $\pi$ rotations.
In summary, our work establishes limits to the temporal resolution reachable by a quantum sensor. We develop our discussion in the framework of the QSL, which is applied to the canonical sensing principle of Ramsey interferometry. We derive expressions for the coherent phase pick-up during a generic sequence consisting of two phase-shifted control pulses and analyze the response regarding time resolution, frequency bandwidth, and sensitivity. We numerically simulate the expected phase response for the single spin of an NV center in diamond, taking the full $S=1$ nature of the spin system and nonlinear driving effects into account. Beyond fundamental aspects in quantum metrology, our work has practical relevance for real-world applications of quantum sensors, such as the mapping of fast magnetization or charge dynamics in materials and integrated devices.
The authors thank John Abendroth, Joseph Renes and Laura Alicia Völker for fruitful discussions. This work has been supported by Swiss National Science Foundation (SNSF) Project Grant No. 200020_212051/1 and by the European Research Council through ERC CoG 817720 (IMAGINE).